OpenStudy (anonymous):

How many real solutions are there to the equation shown below? x2 + 2x + 6 = 0

OpenStudy (anonymous):

to know, you'll have to factor ^_^

OpenStudy (amistre64):

none: the discriminate is negative: 4 -24 = -20

OpenStudy (anonymous):

how come? .-.

OpenStudy (anonymous):

for a quadratic equation ax^2 + bx+ c Disc = b^2 -4ac if D<0 roots are imaginary

OpenStudy (amistre64):

that (b^2 -4ac) is negative

OpenStudy (anonymous):

i qot this we i factored it .

OpenStudy (amistre64):

right, the solutions are imaginary ;)

OpenStudy (anonymous):

i was thinking it Was 2 real solutions .

OpenStudy (amistre64):

if its got an "i" in it; that means "i"maginary. Its a cruel joke really, such a useful complex number and all, but to be dubbed imaginary all your life is kinda a misnomer

OpenStudy (radar):

\[x=-1\pm \sqrt{-5}\]

OpenStudy (amistre64):

it simply means it exists, but not in the realm of the "real" numbers. So it tends to not get invited to all the social gatherings and all the other numbers that hang out with it soon leave ...

OpenStudy (anonymous):

oh, I missed that ^_^"

OpenStudy (anonymous):

so i quess It foold Mee uqqh .

OpenStudy (anonymous):

\[ax^2 + bx + c = 0 \] \[\implies x=\frac{-b\pm \sqrt{b^2-4(a)(c)}}{2(a)}\] \[b^2 - 4(a)(c) < 0 \implies \text{No real solutions}\] Because the thing under the square root would be negative. \[b^2 - 4(a)(c) = 0 \implies \text{One real solution}\] Because the \(\pm\) bit would be \(\pm\) 0. \[b^2 - 4(a)(c) > 0 \implies \text{Two real solutions}\]

OpenStudy (amistre64):

in math, you have to use maths definitions...not english definitions lol

OpenStudy (amistre64):

"real" in math means: not having an "i"

OpenStudy (anonymous):

idk , I dont qet it Maybe one solution

OpenStudy (radar):

\[x=-1\pm \sqrt{5 \times i ^{2}}=-1\pm i \sqrt{5}\]

OpenStudy (amistre64):

No SOLUTIONS

OpenStudy (amistre64):

no REAL solutions lol

OpenStudy (anonymous):

If the part under the square root is negative you have no real soltions.

OpenStudy (anonymous):

but how do yu find if theres Some solutions or not how yu determine

OpenStudy (anonymous):

Look at the quadratic equation.

OpenStudy (anonymous):

Have you learned the quadratic equation?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

Home sweet home:

OpenStudy (anonymous):

Ok. So when you plug in a, b, and c into the quadratic equation. You see the part with the square root?

OpenStudy (anonymous):

What is the domain of the function F(x) graphed below?

OpenStudy (anonymous):

Ohh okayy i qet tht now .

OpenStudy (amistre64):

the set of all Real numbers

OpenStudy (anonymous):

The domain of the function is all the values you can give x and get real solutions for F(x). Since there are no vertical asymptotes or holes in the graph, the domain is \((-\infty,\infty)\)

OpenStudy (anonymous):

amistre64 Wats tht yu wrote the set of all real numbers

OpenStudy (anonymous):

-1 < x < 1

OpenStudy (amistre64):

Yes; the domain of your graph there is the "set of all Real numbers". All the numbers that live on the number line are in the domain

OpenStudy (amistre64):

the domain for f(x) = (x-1)^2 -1 is; any and all real numbers. The graph extends forever to the left, and forever to the right with no restrictions

OpenStudy (anonymous):

What is the y-coordinate of the vertex of a parabola with the following equation? y = x2 - 8x + 18

OpenStudy (amistre64):

(-b/2a, f(-b/2a) :)

OpenStudy (amistre64):

x=4; y = 4^2 -8(4) + 18

OpenStudy (amistre64):

16-32+18 = -16+18 = 2 (4,2)

OpenStudy (anonymous):

itss 18 right

OpenStudy (amistre64):

18 is the "y intercept" that is not the vertex for this graph.

OpenStudy (amistre64):

the vertex for this graph is (4,2) so the y coordinate of the vertex is 2

OpenStudy (amistre64):

the vertex is where the graph bends and starts going back around; its kinda like the "handle" of the graph. we can use it to move the graph anywhere we want of the graph paper

OpenStudy (anonymous):

ohh Okay but I thought it was 18 because i factor it & i thouqht they was asking for tha y intercept But read it to fast & put tha wrong answer . dummie Mee

OpenStudy (amistre64):

lol... I do that alot as well; I read what I want to see in it and get a wrong answer ;)